# Large Language Models as Span Annotators

### v2

@@ 6 Discussion
Span annotation as an evaluation method is not suitable for certain NLG evaluation tasks, such as annotating omissions or rating the overall text style. Although we tried to experiment with a representative sample of models, prompts, and tasks, our choice is necessarily constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their inherent ambiguity, the annotation guidelines, or the qualification level of human annotators.
7 Conclusion
We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
@@ Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@
The human evaluation study was approved by the internal ethics committee of our institution. Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). Annotators were pre-selected on the basis of their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B.2 for details). The domains were selected so that they do not contain sensitive or potentially offensive content. We do not collect demographic data about participants.
@@ A.1 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use the following local models for span annotations:
We also use the following models for generating texts in D2T-Eval:
@@
We run the models on a set of GPUs, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
@@ A.2 Proprietary Models
A.3 Parsing the Output with Reasoning Traces
At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning (ollama models [^15] https://github.com/ollama/ollama/issues/8529). Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
Appendix B D2T-Eval Details
In this section, we provide more details on the process of collecting data (Section B.1) and annotations (Section B.2) for D2T-Eval.
B.1 Collecting Data
To collect a novel dataset for D2T-Eval that is not subject to data contamination (cf. Section 4.4), we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
Domains
The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata. We selected openweather and gsmarena as the most suitable for the data collection process. Additionally, we collect the new football domain from RapidAPI - API-Football. [^16] Its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
Generating model outputs.
To obtain output texts, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 Team et al. (2024) and Phi-3.5 Abdin et al. (2024) (cf. Section A.1). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. We provide the prompts in Appendix D.
Note on annotating models’ own outputs.
Two of the models (Llama 3.3 and GPT-4o) are also used later to collect annotations, which can reveal self-preference bias Zheng et al. (2023); Wataoka et al. (2024). Since we do not rank the quality of generated outputs, we do not consider it a flaw but rather a ground for future investigation. Moreover, we hypothesize that since span annotations are more grounded than direct assessment, the self-preference bias will not be as prominent.
@@ B.2 Collecting annotations
Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the annotations collected for D2T-Eval.
@@
For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020). Eventually, we settled on the list of annotation categories described in Appendix C (Table 5).
Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), also covering coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
Additionally, we agreed upon a list of general recommendations for collecting the annotations:
  • The annotator should always annotate the longest continuous span (i.e., the whole fact instead of a single word).
  • The annotator should annotate according to their own knowledge and only the spans that they are sure about.
  • The annotator should ignore subjective statements such as "a lightweight smartphone".
  • The outputs in openweather can contain both precise numbers (e.g. 10.71°C) and the rounded ones (e.g. 11°C) as long as they agree with the data.
We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 6 and 7).
@@
With the annotation guidelines established, we proceeded to collect our own internally annotated gold data. The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study. Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{iaa}}$, containing 12 examples annotated commonly by each annotator. Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
Qualification task
To gather a pool of skilled annotators, we set up a qualification task on Prolific.com. For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$. For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$ score higher than 0.5 w.r.t. our ground-truth solution to the qualification task. Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline. We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
Annotating test data
We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$. Out of 50 selected participants, 45 participated in the main task (=90% turnover rate). We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (that is, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task). All 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Furthermore, 475 outputs (39.6%) were annotated by an additional annotator. We payed all the annotators an average reward of 10.70 GBP / hour. We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.171717For other experiments, we use only the outputs from the first annotator as the reference data.
B.3 Annotation Interface
We implement our annotation interface using factgenie. Figure 5 shows the main annotation interface, including an example from the openweather domain.
Figure 5: Screenshot of the factgenie annotation interface used for collecting span annotations.
@@ Appendix C Annotation Categories
Tables 5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
@@
Table 7: Annotation categories for the MT-Eval task. The categories are adopted from Da San Martino et al. (2019).
@@ Appendix D Prompts
Figures 6, 7, 8 and 9 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
Figure 10 shows the base prompt we used for MT-Eval, and Figure 11 shows the base prompt we used for Propaganda.
Finally, Figure 12 shows the prompt we used for generating the outputs for D2T-Eval.
@@
Appendix E Results
In this section, we provide detailed results of our experiments.
The results for D2T-Eval are included in Sections E.1, E.2, E.3 and E.4:
The results for MT-Eval are included in Sections E.5 and E.6:
The results for Propaganda are included in Section E.7: Tables 20 and 21.
The confusion matrices are included in Section E.8: Figures 13, 14 and 15.
The manual evaluation results are included in Section E.9: Tables 22 and 23.
E.1 D2T-Eval – Main Results
@@
Table 8: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure 2 for visualizaton of this table.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2981 2.5 28.8 50.3
Llama 3.3 3214 2.7 7.4 65.5
GPT-4o 2284 1.9 4.8 66.3
Claude 3.7 2865 2.4 22.5 57.2
DeepS. R1 1387 1.2 44.2 56.8
o3-mini 1836 1.5 35.6 58.0
Gem. 2-FT 2517 2.1 28.9 54.3
Table 9: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.2 D2T-Evalfootball domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.512 0.288 0.484 0.296 0.498 0.292 0.491 0.199 0.197 0.455
GPT-4o 0.422 0.347 0.502 0.268 0.388 0.302 0.438 0.136 0.190 0.427
Claude 3.7 0.612 0.360 0.511 0.391 0.555 0.375 0.532 0.157 0.254 0.587
DeepS. R1 0.561 0.431 0.670 0.275 0.428 0.336 0.522 0.186 0.276 0.659
o3-mini 0.610 0.470 0.608 0.384 0.496 0.422 0.547 0.125 0.328 0.645
Gem. 2-FT 0.588 0.419 0.617 0.347 0.512 0.379 0.560 0.181 0.307 0.647
Table 10: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1053 2.6 26.8 54.1
Llama 3.3 977 2.4 6.2 71.5
GPT-4o 761 1.9 2.8 72.8
Claude 3.7 1051 2.6 13.8 63.1
DeepS. R1 486 1.2 41.2 67.1
o3-mini 642 1.6 30.5 70.8
Gem. 2-FT 755 1.9 29 61.3
Table 11: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the football domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.3 D2T-Evalgsmarena domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.165 0.107 0.247 0.122 0.282 0.114 0.263 0.149 0.077 0.456
GPT-4o 0.127 0.107 0.269 0.082 0.205 0.093 0.233 0.140 0.062 0.453
Claude 3.7 0.214 0.176 0.348 0.152 0.300 0.163 0.322 0.159 0.140 0.656
DeepS. R1 0.223 0.199 0.405 0.089 0.180 0.123 0.249 0.126 0.160 0.686
o3-mini 0.172 0.225 0.473 0.115 0.241 0.152 0.320 0.168 0.192 0.689
Gem. 2-FT 0.234 0.218 0.426 0.128 0.249 0.161 0.314 0.153 0.161 0.675
Table 12: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 809 2.0 40.2 42.5
Llama 3.3 890 2.2 15 73.9
GPT-4o 670 1.7 10.5 59.2
Claude 3.7 584 1.5 47.2 48.9
DeepS. R1 208 0.5 69.2 53.8
o3-mini 272 0.7 66.2 49.1
Gem. 2-FT 419 1.1 54.2 45.1
Table 13: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the gsmarena domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.4 D2T-Evalopenweather domain
Model $\rho$ Precision Recall F1 $\gamma$ $S_{\emptyset}$
Hard Soft Hard Soft Hard Soft $\Delta$
Llama 3.3 0.209 0.112 0.327 0.118 0.344 0.115 0.335 0.220 0.053 0.272
GPT-4o 0.404 0.196 0.354 0.160 0.289 0.176 0.318 0.142 0.122 0.377
Claude 3.7 0.550 0.274 0.411 0.292 0.437 0.283 0.424 0.141 0.189 0.429
DeepS. R1 0.454 0.241 0.436 0.143 0.258 0.180 0.324 0.144 0.124 0.533
o3-mini 0.552 0.362 0.496 0.267 0.366 0.307 0.422 0.115 0.259 0.481
Gem. 2-FT 0.458 0.226 0.415 0.249 0.458 0.237 0.435 0.198 0.155 0.368
Table 14: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 1119 2.8 19.2 52.2
Llama 3.3 1347 3.4 1 55.7
GPT-4o 853 2.1 1.2 66.2
Claude 3.7 1230 3.1 6.5 56.1
DeepS. R1 693 1.7 22.2 50.5
o3-mini 922 2.3 10 51.8
Gem. 2-FT 1343 3.4 3.5 53.2
Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the openweather domain of D2T-Eval. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.5 MT-Eval – Main Results
@@
Table 16: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
Human 2090 0.7 66.0 14.5
Llama 3.3 6361 2.3 6.2 17.4
GPT-4o 4866 1.7 7.0 15.9
Claude 3.7 3782 1.4 30.6 15.9
DeepS. R1 2586 0.9 36.3 15.1
o3-mini 3039 1.1 35.8 13.8
Gem. 2-FT 2181 0.8 50.0 15.2
Table 17: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.6 MT-Eval – By Language
@@
Table 18: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
Language Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
en-cs Model avg. 600 1.4 27.0 16.6
Human 399 0.7 66.1 13.0
en-es Model avg. 417 1.1 38.9 18.8
Human 248 0.6 70.3 10.3
en-hi Model avg. 396 1.3 26.2 19.0
Human 222 0.5 71.2 10.7
en-is Model avg. 563 1.9 14.3 15.7
Human 752 2.5 18.3 16.6
en-ja Model avg. 471 1.3 28.7 11.1
Human 118 0.2 87.5 14.8
en-ru Model avg. 500 1.3 25.9 18.2
Human 287 0.7 58.7 19.4
en-uk Model avg. 436 1.5 25.4 17.8
Human 208 0.7 64.3 12.3
en-zh Model avg. 420 1.2 34.6 7.2
Human 171 0.2 85.1 6.6
Table 19: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language. Ann=# of annotations, Ann/Ex=annotations per example. Ex. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.7 Propaganda – Main Results
@@
Table 20: IAA between reference and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
Annotator Ann. Ann/Ex Ex. w/o annChars/Ann
@@
Table 21: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda. Ann=# of annotations, Ann/Ex=annotations per example. w/o ann=examples without annotations, Chars/Ann=# characters per annotation.
E.8 Confusion Matrices
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 13: Confusion matrices comparing human annotations (rows) with model predictions (columns) for D2T-Eval. Categories: Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other (cf. Table 5). For the model average, see Figure 3.
(a) Llama 3.3
(b) DeepSeek-R1
(c) GPT-4o
(d) o3-mini
(e) Gemini 2.0 Flash Thinking
(f) Claude 3.7 Sonnet
Figure 14: Confusion matrices comparing human annotations (rows) with model predictions (columns) for MT-Eval. Categories: Major and Minor errors (cf. Table 6).
Figure 15: Confusion matrix of category counts averaged across models for Propaganda (see Table 7 for the description of categories).
E.9 Manual Analysis
Annotation Reason

### v3

@@ 7 Conclusion
Although we aimed to select a representative sample of models, prompts and tasks, our choice is constrained by our limited time frame and budget. Our estimates of the upper-bound IAA for each task are difficult to establish and depend on many factors, such as the chosen annotation categories, their ambiguity, the annotation guidelines, or the qualification level of human annotators. The estimates are also not readily available for existing datasets and require additional data collection.
As an evaluation method, span annotation is not well-suited for certain NLG evaluation tasks such as annotating omissions or rating the overall text style. In these cases, it is best to combine span annotation with other evaluation methods.
@@ Acknowledgments
This work was funded by the European Union (ERC, NG-NLG, 101039303) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
@@
The human evaluation study was approved by the internal ethics committee of our institution. Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). Annotators were pre-selected on the basis of their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research. The domains were selected so that they do not contain sensitive or potentially offensive content. We do not collect demographic data about participants.
@@ A.1 Open Models
We run the local models using the ollama framework in 4-bit quantization. Specifically, we use llama3.3:70b and deepseek-r1:70b for span annotations. We also use gemma2:2b and phi3.5:3.8b for generating texts in D2T-Eval.
@@
We run the models using several GPU variants, including NVIDIA H100 NVL (95G), AMD MI210 (64G), and NVIDIA RTX 3090 (24G).
@@ A.2 Proprietary Models
A.3 Web Interface
We implement our span annotation process using factgenie kasner2024factgenie: a tool that supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
Figure˜12 shows samples of our annotation interface implemented in factgenie for human annotators, including data visualizations from the football and openweather domains.
Appendix B Annotating D2T-Eval
B.1 Generating Outputs
For generating the outputs for the structured inputs we collected, we use two larger models – Llama 3.3 70B grattafiori2024llama and GPT-4o hurst2024gpt – and two smaller models – Gemma 2 2B team2024gemma and Phi-3.5 3.8B abdin2024phi. See more details on the models in Appendix˜A and prompts in Appendix˜D.
@@ B.2 Collecting annotations
For the annotation guidelines, we went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works kasner2024traditional; thomson2020gold. We settled on the following annotation categories (see Table˜5 for details): semantic accuracy errors due to information Contradictory to the input, Not Checkable, or Misleading; any Incoherent and Repetitive content, and any Other errors.
@@
With the annotation guidelines established, we proceeded to collect our own internally annotated gold data: $\mathcal{D}_{\texttt{dev}}$, which contains 84 examples annotated individually by one of 7 annotators (12 examples per annotator) and $\mathcal{D}_{\texttt{iaa}}$, which contains 12 examples annotated commonly by all annotators. [^9] We selected an example for each of the 4 domains and 3 models. The purpose of $\mathcal{D}_{\texttt{dev}}$ is to create a high-quality development set for the model prompting study, while the purpose of $\mathcal{D}_{\texttt{iaa}}$ is to pre-select skilled crowdworkers and quantify the performance of crowdworkers during data collection. Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
Crowdsourcing annotations
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations.
  • Qualification task: For the qualification task, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{iaa}}$. We invited annotators with the $F_{1}$ score higher than 0.5 w.r.t. our internal annotations for the main task.
  • Main task: Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) qualified. Of these, 45 annotators accepted (=90% turnover rate). For annotating the data in $\mathcal{D}_{\texttt{test}}$, we presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{iaa}}$ (that is, the examples that we did not use for the qualification task). All the 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Furthermore, 475 outputs (39.6%) were annotated by an additional annotator. [^10] We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2. For other experiments, we use only the outputs from the first annotator as reference data.
For the qualification task, we paid all the annotators an average reward of 9.58 GBP / hour regardless of the qualification outcome. For the main task, we pay all the annotators an average reward of 10.70 GBP / hour.
@@ Appendix C Annotation Categories
Tables˜5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
@@
Table 7: Annotation categories for the Propaganda task. The categories are adopted from dasanmartino2019finegrained.
@@ Appendix D Prompts
Here, we provide the model prompts:
  • Figures˜5, 6, 7 and 8 show the prompts for the D2T-Eval that we use for the experiments in Section˜5.1.
  • Figure˜9 shows the base prompt we used for MT-Eval.
  • Figure˜10 shows the base prompt we used for Propaganda.
  • Figure˜11 shows the prompt we used for generating the outputs for D2T-Eval.
@@
(a) Data visualization – openweather
(b) Interface for highlighting spans
Figure 12: Samples from the factgenie annotation interface used for collecting span annotations.
Appendix E Examples
In Tables˜8, 9 and 10, we show examples of the annotated outputs for our tasks. Figure˜12 shows our annotation interface.
Source (part of the visualization) [Uncaptioned image]
Annotated by Llama 3 Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaynot check.. Temperatures will remain relatively cool, hovering around freezing point or slightly below. Winds will be light to moderate throughout the period, with some gusts expected on Fridaynot check.. The overall weather pattern suggests a generally cloudy and potentially wet week.
Annotated by o3-mini Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaycont.. Temperatures will remain relatively cool, hovering around freezing point or slightly below. Winds will be light to moderate throughout the period, with some gusts expected on Fridaycont.. The overall weather pattern suggests a generally cloudy and potentially wet weekcont..
Annotated by human Bear Creek will experience mostly overcast skies for the next five days, with a chance of occasional showers on Wednesday and Thursdaycont.. Temperatures will remain relatively cool, hovering around freezing point or slightly below. Winds will be light to moderate throughout the period, with some gusts expected on Fridaymislead.. The overall weather pattern suggests a generally cloudy and potentiallycont. wetcont. weekcont..
Table 8: Example for span annotation of D2T-Eval in the weather domain with contradictorycont, misleadingmislead., and not-checkeablenot check. as error span categories. In the annotation interface, the visualization is interactive. The top part (not visible here) shows the place name and date timeline.
Annotated by Llama 3 “¿Es su cumpleaños pronto, no lo es?”major Ivory le preguntó,minor recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated by o3-mini “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodó, un deslumbramiento tristemajor en sus ojos azules claros.
Annotated by human “¿Es su cumpleaños pronto, no lo es?”minor Ivory le preguntó, recordando que el cumpleaños de la princesa era en unos días. Kari nodómajor, un deslumbramiento triste en sus ojos azules claros.
[…]
Appendix F Results
Here, we provide detailed results of our experiments:
@@ Appendix F Results
Table 11: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. See Figure˜2 for visualizaton of this table.
@@
Table 12: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval – average across languages.
@@
Table 13: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language (average across models).
@@
Table 14: Evaluation of human and LLM annotations using $\mathcal{P}_{\texttt{base}}$ on the Propaganda.
Annotator Ann Ann/Ex w/o% Char/Ann
Human 2981 2.5 28.8 50.3
Llama 3.3 3214 2.7 7.4 65.5
GPT-4o 2284 1.9 4.8 66.3
Claude 3.7 2865 2.4 22.5 57.2
DeepS. R1 1387 1.2 44.2 56.8
o3-mini 1836 1.5 35.6 58.0
Gem. 2-FT 2517 2.1 28.9 54.3
Table 15: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on D2T-Eval. Ann=# of annotations, Ann/Ex=ann. per example. w/o=% ex. without annotations, Char/Ann=# chars per ann.
Annotator Ann. Ann/Ex w/o% Char/Ann
Human 2090 0.7 66.0 14.5
Llama 3.3 6361 2.3 6.2 17.4
GPT-4o 4866 1.7 7.0 15.9
Claude 3.7 3782 1.4 30.6 15.9
DeepS. R1 2586 0.9 36.3 15.1
o3-mini 3039 1.1 35.8 13.8
Gem. 2-FT 2181 0.8 50.0 15.2
Table 16: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on MT-Eval. See Table˜15 for the legend.
Lang. Annot. Ann. Ann/Ex w/o% Char/Ann
en-cs Model 600 1.4 27.0 16.6
Human 399 0.7 66.1 13.0
en-es Model 417 1.1 38.9 18.8
Human 248 0.6 70.3 10.3
en-hi Model 396 1.3 26.2 19.0
Human 222 0.5 71.2 10.7
en-is Model 563 1.9 14.3 15.7
Human 752 2.5 18.3 16.6
en-ja Model 471 1.3 28.7 11.1
Human 118 0.2 87.5 14.8
en-ru Model 500 1.3 25.9 18.2
Human 287 0.7 58.7 19.4
en-uk Model 436 1.5 25.4 17.8
Human 208 0.7 64.3 12.3
en-zh Model 420 1.2 34.6 7.2
Human 171 0.2 85.1 6.6
Table 17: Statistics of models (averaged) and human annotators using $\mathcal{P}_{\texttt{base}}$ on the MT-Eval separately for each language. See Table˜15 for the legend.
Annotator Ann. Ann/Ex w/o% Char/Ann
@@
Table 18: Statistics of models and human annotators using $\mathcal{P}_{\texttt{base}}$ on the Propaganda. See Table˜15 for the legend.
Annotations Explanations